# Local convergence rates for Wasserstein gradient flows and McKean-Vlasov equations with multiple stationary solutions

### v2

@@ 2.5 Auxiliary regularization results
Let us recall two results. The first is proven in [15, Lemma 4.9].
@@
The second result is from [50, Corollary 4.3].
@@
$q_{t}=\left(1+\eta L+\frac{4\eta\lambda}{\sigma^{2}}+\frac{L^{2}\eta^{2}}{2}\right)\left(\frac{\kappa_{1}}{1-e^{-\kappa_{1}t}}+\frac{1}{2}tL^{2}e^{(2\kappa_{1}+4L)t}\right)\,.$ (28)
@@ 2.6 Conclusion
$\partial_{t}|X_{t}-Y_{t}|^{2}=-2(X_{t}-Y_{t})\cdot\left(\nabla E_{\rho_{t}}(X_{t})-\nabla E_{\rho_{*}}(Y_{t})\right)\leqslant(\kappa_{1}+L)|X_{t}-Y_{t}|^{2}+L\mathcal{W}_{2}^{2}\left(\rho_{t},\rho_{*}\right)\,,$
@@ 3.1.1 The associated fixed-point problem
$\mathcal{F}(\rho)-\mathcal{F}(\rho_{\psi_{*}})\leqslant\eta\left(\sigma^{2}+\frac{4\eta\theta\ell^{2}}{(1-\alpha)^{2}}\right)\mathcal{I}\left(\rho|\Gamma(\rho)\right)\,,$
i.e. (NL-LSI) holds over $\mathcal{A}=\{\rho\in\mathcal{P}_{2},\ \varphi(\rho)\in\mathcal{A}^{\prime}\}$.
@@
Under Assumption 6, let $\psi_{*}\in\mathcal{K}^{\prime}$. Assume that $f$ is differentiable at $\psi_{*}$ with $|\nabla_{\mathbb{H}}f(\psi_{*})|<1$. Then there exist $\delta,\overline{\eta}>0$ such that (NL-LSI) holds on $\mathcal{A}=\mathcal{B}_{\mathcal{W}_{2}}(\rho_{\psi_{*}},\delta)$.
@@ 3.2.1 The one-dimensional double-well case
for $m\neq m_{+}$ and $\alpha(m_{+})=f^{\prime}(m_{+})$ is continuous on $\mathbb{R}_{+}$ with values in $(0,1)$ on $(0,\infty)$. Moreover, since $f^{\prime}(m)\rightarrow 0$ as $m\rightarrow\infty$, so does $\alpha$. As a consequence, for any $\varepsilon>0$, $\alpha_{\varepsilon}:=\sup_{m>\varepsilon}\alpha(m)<1$, which proves (48). Proposition 9 then gives (49) for $\rho\in\mathcal{P}_{2}(\mathbb{R})$ with $m(\rho)\geqslant\varepsilon$. The case $m(\rho)\leqslant-\varepsilon$ is obtained by symmetry. ∎
@@
$\partial_{t}m_{t}=-\int_{\mathbb{R}}x^{3}\rho_{t}(\mathrm{d}x)+m_{t}\leqslant-\mathbb{E}(Y_{t}^{3})+m_{t}\,,$
@@
The proof of (52) is then exactly the same as the proof of (46) from (45). ∎
@@ 3.2.2 Localization in the low-temperature regime
It remains to show that Corollary 10 applies. To emphasize the dependency on the temperature, we write $f_{\sigma}$ the function (37). We have to prove that $|\nabla f_{\sigma}(\psi_{*,\sigma})|<1$ for $\sigma$ small enough. From (43) and (44),
@@
Notice that $c$ is independent from $\sigma$, and thus, in Proposition 9, (38) holds on $\mathcal{A}^{\prime}=\mathcal{B}\left(\varphi(a),r\right)$ with $r>0$ and $\alpha\in[0,1)$ independent from $\sigma\in(0,\sigma_{0}^{\prime\prime}]$ for some $\sigma_{0}^{\prime\prime}>0$. If, moroever, $x\mapsto V(x)+|\varphi(x)-\psi|^{2}$ is $k$-strongly convex for some $k>0$ for all $\psi\in\mathcal{A}^{\prime}$, the Bakry-Emery criterion shows that $\Gamma(\rho)$ satisfies an LSI with constant $\eta=k\sigma^{2}$ for all $\rho$ such that $\varphi(\rho)\in\mathcal{A}^{\prime}$, and thus for all $\rho\in\mathcal{A}:=\mathcal{B}_{\mathcal{W}2}(\delta,r_{0})$ provided $r_{0}$ is small enough (independently from $\sigma$). Following the proof of Proposition 9 (which in fact does not require (U-LSI) to hold globally on $\mathcal{P}(\mathbb{R}^{d})$, but only on $\mathcal{A}$), we obtain for $\rho\in\mathcal{A}$ the local non-linear LSI
@@
In the case (1) with $W(x,y)=\theta|x-y|^{2}$, taking $\theta>-\inf\nabla^{2}V/2$ (as in [49, Theorem 2.3]), we get that (53) holds with $\overline{\eta},C$ that are uniform over $\sigma$ small enough in a neighborhood of $\delta_{a}$. At first this seems to be a desirable property for optimization (and in fact this is the idea underlying consensus-based optimization). This should be mitigated by the following observations.
@@ 4 Vlasov-Fokker-Planck equation
Entropic long-time convergence for (62) have been established in some cases (with the energy corresponding to the granular media case (10)) in [40, 27] (using uniform LSI for the associated particle system), [14] (with global non-linear LSI) or [43] (with uniform conditional LSI for the associated particle system and a smallness assumption on the interaction). All these works establish global convergence toward a unique stationary solution and thus do not cover cases with several stationary solutions.
@@
$\partial_{t}\mathcal{F}_{k}(\rho_{t})=-\sigma^{4}\int_{\mathbb{R}^{2d}}\left|\nabla_{v}\ln\frac{\rho_{t}}{\Gamma_{k}(\rho_{t})}\right|^{2}\mathrm{d}\rho_{t}\,.$ (64)
@@
Then, taking $a=(3+4M^{2})^{-1}$ in the definition of $\mathcal{L}$, for all $t\geqslant 0$,
$\partial_{t}\mathcal{L}(\rho_{t})\leqslant-\frac{\sigma^{4}}{4}\mathcal{I}\left(\rho_{t}|\Gamma_{k}(\rho_{t})\right)\,.$
@@
Set $h_{t}=\rho_{t}/\Gamma_{k}(\rho_{t})$. As computed in the proof of [39, Proposition 3] (notice that the friction parameter $\gamma$ of [39] is $1$ in the present case, which besides can always be enforced by a suitable time rescaling, up to rescaling $E_{\rho}$ and $\sigma^{2}$), for any $2d\times 2d$ matrix $A$ of the form
$A=\begin{pmatrix}a_{11}I_{d}&a_{12}I_{d}\\ a_{21}I_{d}&a_{22}I_{d}\end{pmatrix}\,,$
writing $M^{\prime}=M\sqrt{a_{11}^{2}+a_{21}^{2}}$, it holds, for all $r>0$,
@@
$\partial_{t}\int_{\mathbb{R}^{2d}}\left|A\nabla\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\leqslant 2\int_{\mathbb{R}^{2d}}\nabla\ln h_{t}\cdot B\nabla\ln h_{t}\mathrm{d}\rho_{t}$
@@
$B=A^{T}A\begin{pmatrix}M^{\prime}rI_{d}&0\\ -I_{d}&(M^{\prime}r-1)I_{d}\end{pmatrix}+\frac{M^{\prime}}{r}\begin{pmatrix}0&0\\ 0&I_{d}\end{pmatrix}\,.$
Taking $a_{ij}=1/\sqrt{2}$ for all $i,j\in\{1,2\}$ we get that $M^{\prime}=M$ and
$B$ $=$ $\begin{pmatrix}(Mr-1)I_{d}&(Mr-1)I_{d}\\ (Mr-1)I_{d}&\left(Mr-1+\frac{M}{r}\right)I_{d}\end{pmatrix}$
$\preceq$ $\begin{pmatrix}-\frac{1}{4}I_{d}&0\\ 0&\left(\frac{1}{2}+2M^{2}+1\right)I_{d}\end{pmatrix}$
taking $r=1/(2M)$, where $\preceq$ stands for the order over symmetric matrices and we used that $xv\leqslant\frac{1}{4}x^{2}+v^{2}$ to bound the cross term. In other words, we have obtained that
$\partial_{t}\int_{\mathbb{R}^{d}}\left|(\nabla_{x}+\nabla_{v})\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\leqslant-\frac{1}{4}\int_{\mathbb{R}^{d}}\left|\nabla_{x}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}+\left(\frac{3}{2}+2M^{2}\right)\int_{\mathbb{R}^{d}}\left|\nabla_{v}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\,.$
Conclusion follows from (64). ∎
@@
In order to follow the argument that led in the elliptic case to Theorem 8, we need to provide an analogue of Corollary 7 in the present hypoelliptic case. This is done by using the Wasserstein-to-entropy short time regularization of [42, Theorem 4.1] instead of Proposition 6 and the entropy-to-Fisher short time regularization of [14, Proposition 5.5], as we now detail.
The following is from [14, Proposition 5.5] (the proof is based on computations similar to Lemma 17).
@@
The following is a particular case of [42, Theorem 4.1].

### v4

@@ 2.5 Auxiliary regularization results
Let us recall two results. The first is proven in [16, Lemma 4.9].
@@
The second result is from [54, Corollary 4.3].
@@
$q_{t}=\left(1+\eta L+\frac{4\eta\lambda}{\sigma^{2}}+\frac{L^{2}\eta^{2}}{2}\right)\left(\frac{\kappa_{1}}{1-e^{-\kappa_{1}t}}+2tL^{2}e^{(2\kappa_{1}+4L)t}\right)\,.$ (34)
@@ 2.6 Conclusion
The relation between the transport inequality (35) and PL inequalities is studied in more general settings for convex functionals in [8].
@@
$\frac{\mathrm{d}}{\mathrm{d}t}|X_{t}-Y_{t}|^{2}=-2(X_{t}-Y_{t})\cdot\left(\nabla E_{\rho_{t}}(X_{t})-\nabla E_{\rho_{*}}(Y_{t})\right)\leqslant(\kappa_{1}+L)|X_{t}-Y_{t}|^{2}+L\mathcal{W}_{2}^{2}\left(\rho_{t},\rho_{*}\right)\,,$
@@ 3.1.1 The associated fixed-point problem
$\mathcal{F}(\rho)-\mathcal{F}(\rho_{\psi_{*}})\leqslant\eta\left(\sigma^{2}+\frac{4\eta\theta\ell^{2}}{(1-\alpha)^{2}}\right)\mathcal{I}\left(\rho|\Gamma(\rho)\right)\,.$
In particular, provided $\mathcal{F}(\rho_{\psi_{*}})$ is the minimum of $\mathcal{F}$ over $\mathcal{A}=\{\rho\in\mathcal{P}_{2}(\mathbb{R}^{d}),\ \varphi(\rho)\in\mathcal{A}^{\prime}\}$, (NL-LSI) holds over $\mathcal{A}$.
@@
Under Assumption 6, let $\psi_{*}\in\mathcal{K}^{\prime}$ be such that $\rho_{\psi_{*}}$ is a local minimizer of $\mathcal{F}$. Assume that $f$ is differentiable at $\psi_{*}$ with $|\nabla_{\mathbb{H}}f(\psi_{*})|<1$. Then there exist $\delta,\overline{\eta}>0$ such that (NL-LSI) holds on $\mathcal{A}=\mathcal{B}_{\mathcal{W}_{2}}(\rho_{\psi_{*}},\delta)$.
@@ 3.1.2 Decomposition of the free energy
Remark 7.
To sum up, when $\psi_{*}\in\mathcal{K}^{\prime}$ is such that $\nabla^{2}_{\mathbb{H}}R(\psi_{*})$ is negative definite and $\|\nabla_{\mathbb{H}}f\|_{*}<1$, then $\rho_{\psi_{*}}$ is a local minimizer of $\mathcal{F}$ and thus Corollary 10 applies.
@@ 3.2.1 The one-dimensional double-well case
for $m\neq m_{+}$ and $\alpha(m_{+})=f^{\prime}(m_{+})$ is continuous on $\mathbb{R}_{+}$ with values in $(0,1)$ on $(0,\infty)$. Moreover, since $f^{\prime}(m)\rightarrow 0$ as $m\rightarrow\infty$, so does $\alpha$. As a consequence, for any $\varepsilon>0$, $\alpha_{\varepsilon}:=\sup_{m>\varepsilon}\alpha(m)<1$, which proves (54). Using that $\mathcal{F}(\rho_{m_{+}})=\inf\mathcal{F}$, Proposition 9 then gives (55) for $\rho\in\mathcal{P}_{2}(\mathbb{R})$ with $m(\rho)\geqslant\varepsilon$. The case $m(\rho)\leqslant-\varepsilon$ is obtained by symmetry. ∎
@@
$\frac{\mathrm{d}}{\mathrm{d}t}m_{t}=-\int_{\mathbb{R}}x^{3}\rho_{t}(\mathrm{d}x)+m_{t}\leqslant-\mathbb{E}(Y_{t}^{3})+m_{t}\,,$
@@
The proof of (58) is then exactly the same as the proof of (52) from (51) (since $\mathcal{F}(\rho_{m_{*}})=\inf\mathcal{F}$ with $m_{*}=0$). ∎
@@ 3.2.2 Localization in the low-temperature regime
It remains to show that Corollary 10 applies. To emphasize the dependency on the temperature, we write $f_{\sigma}$ the function (43). We have to prove that $|\nabla f_{\sigma}(\psi_{*,\sigma})|<1$ for $\sigma$ small enough (thanks to Remark 7, this will imply that $\rho_{*,\sigma}$ is a local minimizer of $\mathcal{F}$). From (49) and (50),
@@
Notice that $c$ is independent from $\sigma$, and thus, in Proposition 9, (44) holds on $\mathcal{A}^{\prime}=\mathcal{B}\left(\varphi(a),r\right)$ with $r>0$ and $\alpha\in[0,1)$ independent from $\sigma\in(0,\sigma_{0}^{\prime\prime}]$ for some $\sigma_{0}^{\prime\prime}>0$. If, moreover, $x\mapsto V(x)+|\varphi(x)-\psi|^{2}$ is $k$-strongly convex for some $k>0$ for all $\psi\in\mathcal{A}^{\prime}$, the Bakry-Emery criterion shows that $\Gamma(\rho)$ satisfies an LSI with constant $\eta=k\sigma^{2}$ for all $\rho$ such that $\varphi(\rho)\in\mathcal{A}^{\prime}$, and thus for all $\rho\in\mathcal{A}:=\mathcal{B}_{\mathcal{W}2}(\delta,r_{0})$ provided $r_{0}$ is small enough (independently from $\sigma$). Following the proof of Proposition 9 (which in fact does not require (U-LSI) to hold globally on $\mathcal{P}(\mathbb{R}^{d})$, but only on $\mathcal{A}$), we obtain for $\rho\in\mathcal{A}$ the local non-linear LSI
@@
In the case (1) with $W(x,y)=\theta|x-y|^{2}$, taking $\theta>-\inf\nabla^{2}V/2$ (as in [53, Theorem 2.3]), we get that (59) holds with $\overline{\eta},C$ that are uniform over $\sigma$ small enough in a neighborhood of $\delta_{a}$. At first this seems to be a desirable property for optimization (and in fact this is the idea underlying consensus-based optimization). This should be mitigated by the following observations.
@@ 4 Vlasov-Fokker-Planck equation
Entropic long-time convergence rates for (68) have been established in some cases (with the energy corresponding to the granular media case (14)) in [43, 29] (using uniform LSI for the associated particle system), [15] (with global non-linear LSI) or [46] (with uniform conditional LSI for the associated particle system and a smallness assumption on the interaction). All these works establish global convergence toward a unique stationary solution and thus do not cover cases with several stationary solutions. In this latter situation, qualitative convergence to the set of stationary solutions is studied in [25].
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{F}_{k}(\rho_{t})=-\sigma^{4}\int_{\mathbb{R}^{2d}}\left|\nabla_{v}\ln\frac{\rho_{t}}{\Gamma_{k}(\rho_{t})}\right|^{2}\mathrm{d}\rho_{t}\,.$ (70)
@@
Then, taking $a=\frac{1}{2}\left(1+2M^{2}+2\frac{L^{2}}{\sigma^{4}}\right)^{-1}$ in the definition of $\mathcal{L}$, for all $t>0$,
$\frac{\mathrm{d}}{\mathrm{d}t}\mathcal{L}(\rho_{t})\leqslant-\frac{\sigma^{4}a}{2}\mathcal{I}\left(\rho_{t}|\Gamma_{k}(\rho_{t})\right)\,.$
@@
This is a computation from the proof of [15, Theorem 2.1], that we recall now (we follow a presentation similar to the proof of [41, Proposition 3]). For conciseness, write $\hat{\rho}_{t}=\Gamma_{k}(\rho_{t})$, $h_{t}=\rho_{t}/\hat{\rho}_{t}$ and let $A$ be a $2d\times 2d$ matrix of the form
$A=\begin{pmatrix}a_{11}I_{d}&a_{12}I_{d}\\ a_{21}I_{d}&a_{22}I_{d}\end{pmatrix}\,.$
For $t_{1},t_{2}>0$, set
$u(t_{1},t_{2})=\int_{\mathbb{R}^{2d}}\left|A\nabla\ln\frac{\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\right|^{2}\mathrm{d}\rho_{t_{1}},$
with $\nabla=\begin{pmatrix}\nabla_{x}\\ \nabla_{v}\end{pmatrix}$. For $t>0$, we decompose
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{2d}}\left|A\nabla\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}=\partial_{t_{1}}u(t,t)+\partial_{t_{2}}u(t,t)=(\star)+(\star\star)\,.$
Step 1 : bounding ($\star$). Since $\hat{\rho}_{t}$ is the invariant measure of the linear kinetic Fokker-Planck equation with fixed energy equal to $E_{\rho_{t}^{x}}$ with generator
$L_{\rho_{t}}=v\cdot\nabla_{x}-(\nabla E_{\rho_{t}^{x}}+v)\cdot\nabla_{v}+\sigma^{2}\Delta_{v}\,,$
the first part is treated with classical computations (e.g. [40, Lemma 8], and below) leading to
$(\star)\leqslant 2\int_{\mathbb{R}^{2d}}\nabla\ln h_{t}\cdot A^{T}AJ\nabla\ln h_{t}\mathrm{d}\rho_{t}\,,$ (72)
where
$J(x)=\begin{pmatrix}0&\nabla^{2}E_{\rho_{t}^{x}}(x)\\ -I_{d}&-I_{d}\end{pmatrix}=:J_{0}+\begin{pmatrix}0&\nabla^{2}E_{\rho_{t}^{x}}(x)\\ 0&0\end{pmatrix}$
is the Jacobian matrix of the drift of
$L_{\rho_{t}}^{*}=-v\cdot\nabla_{x}+(\nabla E_{\rho_{t}^{x}}-v)\cdot\nabla_{v}+\sigma^{2}\Delta_{v}\,,$
the dual of $L_{\rho_{t}}$ in $L^{2}(\hat{\rho}_{t})$. For the reader’s convenience, let us give a self-contained proof of (72) (keeping computations formal and referring to [15] for technical justifications). We start by some preliminary considerations. First, notice that, for a smooth function $\varphi$ with compact support,
$\int_{\mathbb{R}^{2d}}\varphi\partial_{t}\rho_{t}=\int_{\mathbb{R}^{2d}}L_{\rho_{t}}\varphi\rho_{t}=\int_{\mathbb{R}^{2d}}\varphi L_{\rho_{t}}^{*}\left(\frac{\rho_{t}}{\hat{\rho}_{t}}\right)\hat{\rho}_{t}\,.$
This being true for an arbitrary $\varphi$, this amounts to
$\frac{\partial_{t}\rho_{t}}{\hat{\rho}_{t}}=L_{\rho_{t}}^{*}h_{t}\,,\qquad\text{hence}\qquad\frac{\partial_{t}\rho_{t}}{\rho_{t}}=\frac{L_{\rho_{t}}^{*}h_{t}}{h_{t}}\,.$ (73)
Moreover, denoting for smooth functions $f,g$
$\Gamma(f,g)=\sigma^{2}\nabla_{v}f\cdot\nabla_{v}g\,,$
which for all $t\geqslant 0$ is the carré du champs associated to $L_{\rho_{t}}$ since
$\Gamma(f,g)=\frac{1}{2}\left[L_{\rho_{t}}(fg)-fL_{\rho_{t}}(g)-gL_{\rho_{t}}(f)\right]$
(and similarly if $L_{\rho_{t}}$ is replaced by $L_{\rho_{t}}^{*}$), we see that, for a positive smooth $f$,
$L_{\rho_{t}}^{*}\left(\ln f\right)=\frac{L_{\rho_{t}}^{*}f}{f}-\frac{\Gamma(f,f)}{f^{2}}=\frac{L_{\rho_{t}}^{*}f}{f}-\Gamma(\ln f)\,.$ (74)
Next, using that $\hat{\rho}_{t}$ is invariant by $L_{\rho_{t}}$, it holds $\int_{\mathbb{R}^{2d}}L_{\rho_{t}}\varphi\hat{\rho}_{t}=0$ for all nice $\varphi$. In particular, for any $g$,
$\int_{\mathbb{R}^{2d}}L_{\rho_{t}}(g)\rho_{t}$ $=\int_{\mathbb{R}^{2d}}L_{\rho_{t}}(g)h_{t}\hat{\rho}_{t}$ (75)
$=-\int_{\mathbb{R}^{2d}}\left[gL_{\rho_{t}}(h_{t})+2\Gamma(g,h_{t})\right]\hat{\rho}_{t}\ =\ -\int_{\mathbb{R}^{2d}}\left[h_{t}L_{\rho_{t}}^{*}(g)+2\Gamma(g,h_{t})\right]\hat{\rho}_{t}.$
Now, since for $t_{1},t_{2}>0$,
$\partial_{t_{1}}u(t_{1},t_{2})$ $=\int_{\mathbb{R}^{2d}}\left(\frac{\mathrm{d}}{\mathrm{d}t_{1}}\left|A\nabla\ln\frac{\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\right|^{2}\right)\rho_{t_{1}}+\int_{\mathbb{R}^{2d}}\left|A\nabla\ln\frac{\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\right|^{2}\partial_{t}\rho_{t_{1}}$
$=\int_{\mathbb{R}^{2d}}\left(2A\nabla\ln\frac{\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\cdot A\nabla\frac{\partial_{t}\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\right)\rho_{t_{1}}+\int_{\mathbb{R}^{2d}}L_{\rho_{t_{1}}}\left(\left|A\nabla\ln\frac{\rho_{t_{1}}}{\hat{\rho}_{t_{2}}}\right|^{2}\right)\rho_{t_{1}},$
we get, for $t>0$,
$(\star)=\int_{\mathbb{R}^{2d}}\left(2A\nabla\ln h_{t}\cdot A\nabla\frac{\partial_{t}\rho_{t}}{\hat{\rho}_{t}}\right)\rho_{t}+\int_{\mathbb{R}^{2d}}L_{\rho_{t}}\left(\left|A\nabla\ln h_{t}\right|^{2}\right)\rho_{t}.$
Therefore, writing $\Phi=|A\nabla\ln h_{t}|^{2}$, using (73) and the equality (75) with $g=\Phi$, and then (74),
(($\star$)( ⋆ ) $=\int_{\mathbb{R}^{2d}}2A\nabla\ln h_{t}\cdot A\nabla\left(\frac{L_{\rho_{t}}^{*}h_{t}}{h_{t}}\right)\rho_{t}-\int_{\mathbb{R}^{2d}}\left[h_{t}L_{\rho_{t}}^{*}(\Phi)+2\Gamma(\Phi,h_{t})\right]\hat{\rho}_{t}$
$=\int_{\mathbb{R}^{2d}}2A\nabla\ln h_{t}\cdot A\nabla\left(L_{\rho_{t}}^{*}(\ln h_{t})+\Gamma(\ln h_{t})\right)\rho_{t}-\int_{\mathbb{R}^{2d}}\left[h_{t}L_{\rho_{t}}^{*}(\Phi)+2\Gamma(\Phi,h_{t})\right]\hat{\rho}_{t}.$ (76)
Focusing on the first term, and noticing that
$\nabla L_{\rho_{t}}^{*}f=L_{\rho_{t}}^{*}\nabla f+J\nabla f\,,$
we see that, denoting by $(B)$ the right-hand side of (72)
$\int_{\mathbb{R}^{2d}}2A\nabla\ln h_{t}\cdot A\nabla L_{\rho_{t}}^{*}(\ln h_{t})\rho_{t}$ $=$ $(B)+\int_{\mathbb{R}^{2d}}2A\nabla\ln h_{t}\cdot L_{\rho_{t}}^{*}\left(A\nabla\ln h_{t}\right)\rho_{t}$
$=$ $(B)+\int_{\mathbb{R}^{2d}}\left(L_{\rho_{t}}^{*}(\Phi)-2\Gamma(A\nabla\ln h_{t})\right)\rho_{t}\,.$
(In the first line, $L_{\rho_{t}}^{*}\left(A\nabla\ln h_{t}\right)$ is understood as applying coordinate-wise $L_{\rho_{t}}^{*}$ to $A\nabla\ln h_{t}$, and similarly $\Gamma$ in the second line is the sum over the coordinates of $\Gamma$ applied coordinate-wise). Notice that the first term of this integral cancels out with the first term of the second integral of (76). Moreover, by chain rule, for $i\in\llbracket 1,d\rrbracket$,
$A\nabla\ln h_{t}\cdot A\nabla\left(|\partial_{v_{i}}\ln h_{t}|^{2}\right)=2\partial_{v_{i}}\ln h_{t}A\nabla\ln h_{t}\cdot A\nabla\partial_{v_{i}}\ln h_{t}=\partial_{v_{i}}\ln h_{t}\partial_{v_{i}}(\Phi)\,.$
Summing over $i$ and multiplying by $\sigma^{2}$ reads
$A\nabla\ln h_{t}\cdot A\nabla\left(\Gamma\left(\ln h_{t}\right)\right)=\Gamma\left(\ln h_{t},\Phi\right)=\frac{\Gamma\left(h_{t},\Phi\right)}{h_{t}}\,.$
This means that the second terms of the two integrals in (76) cancel out. We have obtained that
$(\star)=(B)-2\int_{\mathbb{R}^{2d}}\Gamma(A\nabla\ln h_{t})\rho_{t}\,.$
The integral being non-positive, this concludes the proof of (72).
Writing $a_{1}=\sqrt{a_{11}^{2}+a_{21}^{2}}$, we bound
(($\star$)( ⋆ ) $\leqslant$ $2\int_{\mathbb{R}^{2d}}\left[\nabla\ln h_{t}\cdot A^{T}AJ_{0}\nabla\ln h_{t}+a_{1}M\left|A\nabla\ln h_{t}\right|\left|\nabla_{v}\ln h_{t}\right|\right]\mathrm{d}\rho_{t}$
$\leqslant$ $\int_{\mathbb{R}^{2d}}\left[\nabla\ln h_{t}\cdot A^{T}A\left(2J_{0}+rI_{2d}\right)\nabla\ln h_{t}+\frac{a_{1}^{2}M^{2}}{r}\left|\nabla_{v}\ln h_{t}\right|^{2}\right]\mathrm{d}\rho_{t}\,,$
for any $r>0$.
[…]
$=$ $2a_{1}\int_{\mathbb{R}^{2d}}\left|A\nabla\ln h_{t}\right|\left|\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{x}\ln\hat{\rho}_{t}\right|\mathrm{d}\rho_{t}$
$\leqslant$ $r^{\prime}\int_{\mathbb{R}^{2d}}\left|A\nabla\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}+\frac{a_{1}^{2}}{r^{\prime}}\int_{\mathbb{R}^{2d}}\left|\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{x}\ln\hat{\rho}_{t}\right|^{2}\mathrm{d}\rho_{t}\,,$
for any $r^{\prime}>0$. By the explicit expression of $\hat{\rho}_{t}$ and (Lip), for $t,s\geqslant 0$,
$\left|\nabla_{x}\ln\hat{\rho}_{t}-\nabla_{x}\ln\hat{\rho}_{t+s}\right|=\frac{1}{\sigma^{2}}\left|\nabla_{x}E_{\rho_{t}^{x}}-\nabla_{x}E_{\rho_{t+s}^{x}}\right|\leqslant\frac{L}{\sigma^{2}}\mathcal{W}_{2}\left(\rho_{t}^{x},\rho_{t+s}^{x}\right)\,.$
Using the continuity equation on $\rho_{t}^{x}$ obtained by integrating (68) over the variable $v$ and using the Benamou-Brenier formulation of the Wasserstein distance [7], as detailed in the proof of [15, Theorem 2.1], we get that
$\limsup_{s\rightarrow 0}s^{-1}\mathcal{W}_{2}^{2}\left(\rho_{t}^{x},\rho_{t+s}^{x}\right)\leqslant\int_{\mathbb{R}^{2d}}\left|\nabla_{v}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\,,$
from which
$\left|\frac{\mathrm{d}}{\mathrm{d}t}\nabla_{x}\ln\hat{\rho}_{t}\right|^{2}\leqslant\frac{L^{2}}{\sigma^{4}}\int_{\mathbb{R}^{2d}}\left|\nabla_{v}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\,.$
[…]
$\partial_{t}\int_{\mathbb{R}^{2d}}\left|A\nabla\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\leqslant\int_{\mathbb{R}^{2d}}\nabla\ln h_{t}\cdot B\nabla\ln h_{t}\mathrm{d}\rho_{t}$
@@
$B=A^{T}A\begin{pmatrix}(r+r^{\prime})I_{d}&0\\ -2I_{d}&(r+r^{\prime}-2)I_{d}\end{pmatrix}+a_{1}^{2}\left(\frac{M^{2}}{r}+\frac{L^{2}}{r^{\prime}\sigma^{4}}\right)\begin{pmatrix}0&0\\ 0&I_{d}\end{pmatrix}\,.$
Taking $a_{ij}=1/\sqrt{2}$ for all $i,j\in\{1,2\}$ we get that
$B$ $=$ $\begin{pmatrix}(r+r^{\prime}-2)I_{d}&(r+r^{\prime}-2)I_{d}\\ (r+r^{\prime}-2)I_{d}&\left(r+r^{\prime}-2+\frac{M^{2}}{r}+\frac{L^{2}}{r^{\prime}\sigma^{4}}\right)I_{d}\end{pmatrix}$
$\preceq$ $\begin{pmatrix}-\frac{1}{2}I_{d}&0\\ 0&\left(1+2M^{2}+2\frac{L^{2}}{\sigma^{4}}\right)I_{d}\end{pmatrix}$
taking $r=r^{\prime}=1/2$, where $\preceq$ stands for the order over symmetric matrices and we used that $-2xv\leqslant\frac{1}{2}x^{2}+2v^{2}$ to bound the cross term. In other words, we have obtained that
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}^{d}}\left|(\nabla_{x}+\nabla_{v})\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\leqslant-\frac{1}{2}\int_{\mathbb{R}^{d}}\left|\nabla_{x}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\\ +\left(1+2M^{2}+2\frac{L^{2}}{\sigma^{4}}\right)\int_{\mathbb{R}^{d}}\left|\nabla_{v}\ln h_{t}\right|^{2}\mathrm{d}\rho_{t}\,.$
Conclusion follows from (70) (the definition of $a$ ensuring that the last term above is compensated by half the dissipation term of (70)). ∎
@@
In order to follow the argument that led in the elliptic case to Theorem 8, we need to provide an analogue of Corollary 7 in the present hypoelliptic case. This is done by using the Wasserstein-to-entropy short time regularization of [45, Theorem 4.1] instead of Proposition 6 and the entropy-to-Fisher short time regularization of [15, Proposition 5.5], as we now detail.
The following is from [15, Proposition 5.5] (the proof is based on computations similar to Lemma 17).
@@
The following is a particular case of [45, Theorem 4.1].